# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

### v2

@@ IV-D Practicality with Consumer-Grade Channels
Despite advancements in brainwave sensors in recent years, consumer-grade devices still have a limited number of sensors. Therefore, we selected three devices to evaluate performance by focusing exclusively on the electrodes specific to each device. The selected devices include the Emotiv Epoc X, which features 14888AF3, F7, F3, FC5, T7, P7, O1, O2, P8, T8, FC6, F4, F8, AF4 electrodes and costs 999 dollars, and the Muse 2, equipped with 4999TP9, AF7, AF8, TP10 electrodes and available at 269.99 euros. Both are budget-friendly options for consumer use. Additionally, we considered the DSI-VR300101010FCz, Pz, P3, P4, PO7, PO8, Oz, which includes 7 electrodes and costs 10,000 dollars. This device, though expensive, has the potential to become more affordable in the future and could be integrated into XR devices.
We selected the same electrodes as used in the devices and trained a model for each device. Table IX represents the results with a lower number of channels. The baseline EER increases from 10.28% to 13.37%, 16.61%, and 18.18% for Emotiv Epoc X, DSI-VR300, and Muse 2, respectively. The reduction from 93 channels in our main dataset to 14, 7, and 4 channels suggests redundancy due to correlated EEG signals across electrodes. Moreover, using averaging techniques and multiple enrollment sessions can further improve the results.
@@ IV-E1 Comparison with Brainwave Authentications Systems
Table VI summarizes the state-of-the-art in multi-session brainwave authentication studies.
@@
TABLE X: Performance metrics for different scenarios with detailed enrollment and verification session information. (Verification samples refer to the number of samples from the same session used in a single verification attempt.)
@@
First All Others 1 10.28±6.3010.28$\pm$ 6.3010.28 ± 6.30 43.33±24.4643.33$\pm$ 24.4643.33 ± 24.46 66.98±25.3066.98$\pm$ 25.3066.98 ± 25.30 80.65±21.4580.65$\pm$ 21.4580.65 ± 21.45
First All Others 4 6.22±5.646.22$\pm$ 5.646.22 ± 5.64 27.86±25.7527.86$\pm$ 25.7527.86 ± 25.75 50.39±30.4050.39$\pm$ 30.4050.39 ± 30.40 64.75±30.0364.75$\pm$ 30.0364.75 ± 30.03
First Second 1 6.87±7.616.87$\pm$ 7.616.87 ± 7.61 32.15±32.2932.15$\pm$ 32.2932.15 ± 32.29 54.04±35.6654.04$\pm$ 35.6654.04 ± 35.66 65.07±34.5265.07$\pm$ 34.5265.07 ± 34.52
First Second 4 3.04±6.233.04$\pm$ 6.233.04 ± 6.23 20.71±33.9820.71$\pm$ 33.9820.71 ± 33.98 33.44±40.5833.44$\pm$ 40.5833.44 ± 40.58 41.26±42.1641.26$\pm$ 42.1641.26 ± 42.16
First Two Third 1 4.87±5.364.87$\pm$ 5.364.87 ± 5.36 23.98±28.2923.98$\pm$ 28.2923.98 ± 28.29 43.59±35.0443.59$\pm$ 35.0443.59 ± 35.04 56.39±35.1556.39$\pm$ 35.1556.39 ± 35.15
First Two Third 4 1.54±2.741.54$\pm$ 2.741.54 ± 2.74 12.24±26.9912.24$\pm$ 26.9912.24 ± 26.99 24.28±35.3324.28$\pm$ 35.3324.28 ± 35.33 32.11±38.7932.11$\pm$ 38.7932.11 ± 38.79
First Two Third 16 0.68±1.750.68$\pm$ 1.750.68 ± 1.75 9.60±26.679.60$\pm$ 26.679.60 ± 26.67 16.08±33.3716.08$\pm$ 33.3716.08 ± 33.37 19.99±36.3119.99$\pm$ 36.3119.99 ± 36.31
@@ IV-E2 Comparison with Biometric Industry Standards
[…]
While our experiments show that brainwave authentication does not yet meet international standards, our previous experiment on the impact of headsets on performance (Sec.IV-B4) revealed that the headset with a higher number of subjects yielded a lower error rate, and we also know that biometrics that meet these standards typically employ datasets with thousands of subjects[53]. Therefore, we conducted an experiment to explore the relationship between the number of subjects used for training the feature extractor and the error rate. The results in Figure 8 indicate a linear relationship between the logarithm of the number of subjects and the EER. It is obvious that the linear decline eventually slows down. The current trend suggests that with a dataset of a few thousand subjects (from 2,000 to 16,000, depending on the device), it may be possible to approach biometric performance standards. This is reasonable when compared to the data volumes typically used in modalities like face recognition.
@@ V Limitations
Our study is subject to certain limitations. First, our stimuli were limited to a general text-based memorization task. Ideally, multiple task types could be included to allow for performance comparisons and identification of the best-performing task. Second, we simulated consumer-grade devices by selecting the same number and channel locations as those found in such devices. However, we did not have a measure to directly compare the quality of sensors. Although sensor quality will improve, the main challenge will remain their number and placement for effective use. It is worth noting that all other multi-session studies in the literature [16, 13, 11, 14] also employ medical-grade data. Finally, we only had access to a single dataset including a notable number of sessions and subjects. However, best practices in state-of-the-art face recognition research recommend training on multiple large datasets and testing on an independent dataset to ensure a more robust evaluation [54]. To simulate this, we split our dataset into a subject-independent set for training feature extraction and evaluation. Moreover, we partitioned the dataset based on device, showing that cross-device evaluation significantly increases error (Tables III and IV.) Achieving robust performance requires large amounts of data to extract distinctive features and account for subject-specific, environmental, and hardware-related variations. For example, in the case of face recognition, Kim et al. [54] used 3 datasets comprising over 350,000 subjects and more than 14.9 million images to train a generalized model. Such large-scale datasets are not yet available in brainwave authentication.
@@ VI Conclusion and Future Work
For future work, there is an urgent need for a public multi-session dataset with a few thousand subjects to ensure brainwave authentication research meets international standards. This dataset should be collected using consumer-grade devices in multiple sessions, and incorporating multiple devices in the same session would be valuable for studying the cross-headset authentication challenge. On the technical side, further research should focus on developing deep learning architectures that better capture the unique features of brainwave data, as well as effective signal preprocessing approaches to remove noise. Adopting an open-source approach would enable researchers to build on each other’s work, promoting reproducibility and collaboration. Further investigation of more advanced attacks on brainwave authentication is necessary.

### v3

@@ IV-D1 Consumer-grade headsets
Despite advancements in brainwave sensors in recent years, consumer-grade devices still have a limited number of sensors. Therefore, we selected three devices to evaluate performance by focusing exclusively on the electrodes specific to each device. The selected devices include the Emotiv Epoc X, which features 14101010AF3, F7, F3, FC5, T7, P7, O1, O2, P8, T8, FC6, F4, F8, AF4 electrodes and costs 999 dollars, and the Muse 2, equipped with 4111111TP9, AF7, AF8, TP10 electrodes and available at 269.99 euros. Both are budget-friendly options for consumer use. Additionally, we considered the DSI-VR300121212FCz, Pz, P3, P4, PO7, PO8, Oz, which includes 7 electrodes and costs 10,000 dollars. This device, though expensive, has the potential to become more affordable in the future and could be integrated into XR devices. We simulated consumer-grade devices by selecting, from the PEERS dataset, the electrodes corresponding to each device’s channel layout. We subsequently trained a separate model for each configuration, following the same subject-disjoint split (230 subjects for training, 100 for evaluation). Table VIII represents the results with a lower number of channels. The baseline EER increases from 10.28% to 13.37%, 16.61%, and 18.18% for Emotiv Epoc X, DSI-VR300, and Muse 2, respectively. The reduction from 93 channels in our main dataset to 14, 7, and 4 channels suggests redundancy due to correlated EEG signals across electrodes. Moreover, using averaging techniques and multiple enrollment sessions can further improve the results.
IV-D2 Regional channel subsets
To analyze which scalp regions contribute most to authentication performance under a fixed channel budget, we grouped electrodes into frontal, central, temporal, and parietal–occipital regions according to their 10–20 labels. We first counted, in each region, how many such symmetric pairs were available and found that the smallest region contained six pairs; we therefore fixed the subset size to six pairs (12 channels) in all regions so that every regional model operated with the same number and type of electrodes. Separate models were trained on each regional subset using the same architecture and protocol as in the full-channel setting, and the resulting EERs are reported in Table IX. The parietal–occipital subset achieved the lowest error rates, followed by frontal and central subsets, while temporal channels yielded slightly higher EERs. This indicates that placing electrodes over parietal–occipital areas preserves authentication performance best, in line with previous EEG biometric studies[35, 31].
Overall, these channel-reduction experiments indicate that a subset of electrodes can retain much of the authentication performance while improving usability. Our consumer-grade headset configurations show that a substantial fraction of the baseline performance can be preserved even with only 4–14 channels, suggesting that low-channel-count devices can be effective when properly configured. Regarding electrode placement, parietal–occipital sites demonstrate the best performance, which may be partly explained by the fact that they are less affected by eye-movement artifacts that mainly disturb electrodes on the forehead close to the eyes [47]. Therefore, these sites are particularly attractive for future consumer-grade devices.Moreover, as discussed in Sections IV-B3 and IV-B4, hardware characteristics affect performance, but this can be mitigated by providing sufficient data from target devices and incorporating such data into model training, which will be crucial for achieving reliable results. Finally, if developers adopt a common set of electrode locations, this would support data pooling across devices and studies, benefiting both authentication and other EEG-based applications.
@@ IV-E1 Comparison with Brainwave Authentications Systems
Table X summarizes the state-of-the-art in multi-session brainwave authentication studies.
@@
TABLE XI: Performance metrics for different scenarios with detailed enrollment and verification session information. (Verification samples refer to the number of samples from the same session used in a single verification attempt.)
@@
First All Others 1 $10.28\pm 6.30$ $43.33\pm 24.46$ $66.98\pm 25.30$ $80.65\pm 21.45$
First All Others 4 $6.22\pm 5.64$ $27.86\pm 25.75$ $50.39\pm 30.40$ $64.75\pm 30.03$
First Second 1 $6.87\pm 7.61$ $32.15\pm 32.29$ $54.04\pm 35.66$ $65.07\pm 34.52$
First Second 4 $3.04\pm 6.23$ $20.71\pm 33.98$ $33.44\pm 40.58$ $41.26\pm 42.16$
First Two Third 1 $4.87\pm 5.36$ $23.98\pm 28.29$ $43.59\pm 35.04$ $56.39\pm 35.15$
First Two Third 4 $1.54\pm 2.74$ $12.24\pm 26.99$ $24.28\pm 35.33$ $32.11\pm 38.79$
First Two Third 16 $0.68\pm 1.75$ $9.60\pm 26.67$ $16.08\pm 33.37$ $19.99\pm 36.31$
@@ IV-E2 Comparison with Biometric Industry Standards
Brainwave authentication is in its early stages, making it crucial to assess progress and identify gaps for real-world implementation. For a biometric system to be deployed in practical applications, it must adhere to international industrial biometric standards. These standards emphasize a very low FAR to ensure security, while maintaining a reasonable FRR to keep the system usable. According to guidelines from NIST (2023)/ISO [^13] https://pages.nist.gov/800-63-3/sp800-63b.htmland the European Border Guard Agency Frontex [^14] https://www.frontex.europa.eu/assets/Publications/Research/Best_Practice_Technical_Guidelines_ABC.pdf, biometric systems should operate at FAR $\leq 0.1\%$ (1 in 1,000). Stricter recommendations from FIDO [^15] https://fidoalliance.org/specs/biometric/requirements/Biometrics-Requirements-v4.0.1-fd-20240522.pdf and more recent NIST (August 2024)/ISO 2024 standards [^16] https://pages.nist.gov/800-63-4/sp800-63b.html suggest a FAR $\leq 0.01\%$ (1 in 10,000). For usability, FRR standards propose a maximum FRR of $\leq 5\%$, allowing at least 19 successful logins out of 20 attempts for genuine users. Table XI presents the FRR at low FAR of 1%, 0.1%, and 0.01%. The results show higher FRR. For instance, using the first session as enrollment and all others as verification, with a duration of 4 seconds, the FRR is 50.39% at FAR 0.1% and 64.75% at FAR 0.01%. This means approximately 9 and 7 successful logins out of 20 attempts for legitimate users, respectively.
While our experiments show that brainwave authentication does not yet meet international standards, our previous experiment on the impact of headsets on performance (Sec.9) revealed that the headset with a higher number of subjects yielded a lower error rate, and we also know that biometrics that meet these standards typically employ datasets with thousands of subjects[16]. Therefore, we conducted an experiment to explore the relationship between the number of subjects used for training the feature extractor and the error rate. The results in Figure 8 indicate a linear relationship between the logarithm of the number of subjects and the EER. It is obvious that the linear decline eventually slows down. The current trend suggests that with a dataset of a few thousand subjects (from 2,000 to 16,000, depending on the device), it may be possible to approach biometric performance standards. This is reasonable when compared to the data volumes typically used in modalities like face recognition.
@@ V Limitations
Our study is subject to certain limitations at three levels: the authentication task, the dataset size, and the recoding hardware.
At the authentication task level, our stimuli were limited to a general text-based memorization task, which imposes non-negligible cognitive load, requires explicit focused attention during authentication, and may limit usability in practical settings; therefore, the usability of this specific task requires further investigation, despite existing general usability studies on brainwave authentication [11]. Ideally, multiple task types could be included to allow for performance comparisons and identification of the best-performing task. Additionally, we did not explicitly address replay attacks or EEG-based liveness detection. These aspects can be mitigated by adapting the approach of Lin et al. [33], which involves varying the stimulus and detecting corresponding changes in EEG responses.
At the dataset size level, we only had access to a single dataset including a notable number of sessions and subjects. However, best practices in state-of-the-art face recognition research recommend training on multiple large datasets and testing on an independent dataset to ensure a more robust evaluation [29]. To simulate this, we split our dataset into a subject-independent set for training feature extraction and evaluation. Moreover, we partitioned the dataset based on device, showing that cross-device evaluation significantly increases error (Tables III and IV.) Achieving robust performance requires large amounts of data to extract distinctive features and account for subject-specific, environmental, and hardware-related variations. For example, in the case of face recognition, Kim et al. [29] used 3 datasets comprising over 350,000 subjects and more than 14.9 million images to train a generalized model. Such large-scale datasets are not yet available in brainwave authentication.
At the hardware level, we simulated consumer-grade devices by selecting the same number and channel locations as those found in such devices. However, we did not have a measure to directly compare the quality of sensors. Although sensor quality will improve, the main challenge will remain their number and placement for effective use. It is worth noting that all other multi-session studies in the literature [34, 8, 59, 6] also employ medical-grade data.
@@ VI Conclusion and Future Work
For future work, there is an urgent need for a public multi-session dataset with a few thousand subjects to ensure brainwave authentication research meets international standards. This dataset should be collected using consumer-grade devices in multiple sessions, and incorporating multiple devices in the same session would be valuable for studying the cross-headset authentication challenge, and across different task paradigms (e.g., passive viewing, motor tasks) to study task similarity. On the technical side, further research should focus on developing deep learning architectures that better capture the unique features of brainwave data, as well as effective signal preprocessing approaches to remove noise. Specifically, the new trend of utilizing attention mechanisms and specialized transformer architectures [42], which have shown promising results in other biometric modalities (e.g., enhanced keystroke dynamics [52]), needs future investigation to see how they can be effectively integrated into EEG-based systems for improved feature extraction and long-term performance. Finally, adopting an open-source approach would enable researchers to build on each other’s work, promoting reproducibility and collaboration. Further investigation of more advanced attacks on brainwave authentication is necessary.
@@ Acknowledgment
Appendix A Session Count Distribution
Figure 9: Histogram of session counts per subject.
Figure 9 shows the session count distribution in our dataset.
Appendix B Architectural Details of Deep Learning Feature Encoders
This appendix provides detailed specifications for the Convolutional Neural Network (CNN) architectures employed as feature encoders in our benchmarking study. In our experiments, CNN-based models trained with Supervised Contrastive Loss (SupConLoss) [28] generally achieved the best performance among the evaluated feature extraction methods, with the strongest results obtained using ResNet1D.
B-A Best-Performing Architecture: ResNet1D
The core concept of the Residual Network was introduced by He et al. [19] to address the degradation problem in very deep neural networks. By incorporating skip connections (residual blocks), ResNet allows information to bypass one or more layers, enabling the stable training of networks with significantly greater depth. In our work, we adopt the ResNet1D backbone proposed by Zheng et al. [62] for EEG anomaly detection, which is based on a ResNet-34-style architecture with temporally oriented convolutional kernels for multi-channel EEG segments.
Architecture Detail
ResNet1D is a deep CNN built from residual blocks that operate along the temporal dimension of the EEG. An initial temporal convolutional block is followed by several stages of stacked residual blocks. Each residual block contains one or more 1D convolutional layers with batch normalization (BN) and rectified linear unit (ReLU) activations, together with a skip connection from input to output. After the last residual stage, global average pooling over time is applied and the resulting feature vector is passed through a fully connected layer, which provides the final feature embedding for comparison.
B-B ShallowConvNet
Introduced by Schirrmeister et al. [49], the ShallowConvNet was designed to closely mimic the successful Filter Bank Common Spatial Pattern (FBCSP) pipeline [2] in an end-to-end deep learning framework. The goal was to demonstrate that a simple deep-learning structure could perform as well.
Architecture Detail
This architecture consists of two convolutional layers, specifically engineered to extract frequency and spatial information, followed by operations that implement log-variance features:
  1. 1.
    Temporal filtering (Block 1): A convolution with kernel size $1\times 25$ is applied along the temporal dimension. This layer learns temporal filters independently across each EEG channel.
  2. 2.
    Spatial filtering (Block 2): A convolution with kernel size $Channels\times 1$ is used to linearly combine the temporal features across electrodes, followed by batch normalization (BN).
  3. 3.
    Feature extraction: The output of the spatial filter then passes through a squaring operation ($x\mapsto x^{2}$), followed by mean pooling (size $1\times 75$, stride $15$), and a logarithmic activation ($\log(x)$). These operations directly implement the log-variance feature computation analogous to FBCSP.
B-C DeepConvNet
Also introduced by Schirrmeister et al. [49], DeepConvNet was designed as a more generic CNN architecture, closer to models used in computer vision such as VGG [54]. The motivation was to explore whether less specialized, deeper networks could capture a wider variety of abstract features from the raw EEG signal, rather than being biased towards the amplitude-based features commonly used in specific task decoding.
Architecture Detail
The network is constructed with four sequential blocks, using smaller kernels and max pooling for feature compression:
  1. 1.
    Block 1: Temporal convolution (size $1\times 10$) followed by a spatial convolution (size $Channels\times 1$), an exponential linear unit (ELU) non-linearity, and max pooling (size $1\times 3$, stride $3$).
  2. 2.
    Blocks 2, 3, 4: These three blocks are stacked consecutively. Each consists of a convolution (size $1\times 10$), an ELU activation, and max pooling (size $1\times 3$, stride $3$). These blocks use small convolutional kernels to extract progressively abstract and compressed features.
  3. 3.
    Output: The final output is flattened and passed through a dense (fully connected) layer, which provides the feature embedding for comparison.
B-D EEGNet
Lawhern et al. [32] introduced EEGNet to create a compact and generalized CNN that could perform well across multiple brain-computer interface (BCI) paradigms (e.g., P300, SMR, ERN) and datasets. The primary design constraint was to achieve this with a minimal number of trainable parameters, making the model suitable for deployment in resource-constrained or real-time BCI systems.
Architecture Detail
EEGNet achieves its efficiency through the factorization of standard convolutions into depthwise and separable convolutions:
  1. 1.
    Block 1 (temporal–spatial filter): This block consists of a temporal convolution followed by a depthwise convolution. Temporal filters learn frequency features. The depthwise filters then apply a unique spatial filter to the output of each temporal filter, extracting frequency-specific spatial patterns.
  2. 2.
    Block 2 (feature mixing): This block employs a separable convolution, which is the combination of a depthwise convolution (providing a temporal summary for each feature map) and a pointwise convolution (a $1\times 1$ convolution). The pointwise convolution acts as an efficient final layer, learning how to optimally combine or mix the feature maps across channels.
[Uncaptioned image] Matin Fallahi is currently pursuing the Ph.D. degree in computer science with the KASTEL – Institute of Information Security and Dependability, Karlsruhe Institute of Technology (KIT), Karlsruhe, Germany. The research focuses on biometric authentication, in particular brainwave-based user authentication using electroencephalography (EEG). This includes the security and robustness of EEG-based biometrics, usability and continuous authentication in extended reality (XR) scenarios, and privacy-preserving processing of biometric data. Fallahi has coauthored several publications on brainwave-based biometrics and related topics in usable security and privacy.
[Uncaptioned image] Patricia Arias-Cabarcos is a Professor of IT Security at the University of Paderborn, where she leads the Human-centered IT Security (HITS) group. Her research focuses on usable authentication, behavioral data privacy, data transparency, and related topics in usable security and privacy. Before joining Paderborn, she was a Senior Researcher at the Karlsruhe Institute of Technology (KIT) from 2019 to 2021, a Humboldt Fellow at the University of Mannheim between 2017 and 2019, and an Assistant Professor at Universidad Carlos III de Madrid (2013–2018), where she also obtained her Ph.D. in Telematic Engineering. She has additionally held positions as a visiting postdoctoral researcher at TU Darmstadt and as a research intern at NEC Laboratories Europe in Heidelberg.
[Uncaptioned image] Thorsten Strufe is Professor of IT Security at KIT, adjunct professor for Privacy and Network Security at TU Dresden, Deputy Speaker of the Cluster of Excellence CeTI (Centre for Tactile Internet with Human-in-the-Loop), and PI in the German Competence Center for IT Security KASTEL, the Research Training Group RoSI (TU Dresden), and the 5G-Lab Germany. Prior to his position at KIT, he was full professor of privacy and security at TU Dresden from 2014 to 2019, received an offer for the Chair of Security and Privacy at University of Freiburg, was assistant professor for Peer-to-Peer Networks at TU Darmstadt from 2009 to 2014, and deputy professor for Secure Distributed Systems at the University of Mannheim throughout 2011. During this time, he was PI in several research groups and collaborative projects, such as the SFB 912 HAEC (Highly Adaptive Energy-Efficient Communication, TU Dresden), the SFB 1053 MAKI (Multi-Mechanism Adaptation for the Future Internet, TU Darmstadt), the research group 733 QuaP2P (Improving the Quality of Peer-to-Peer Systems, TU Darmstadt) and various collaborative projects within the funding programmes of EU and BMBF.